C = 50 + 0.80 (Y – T) I = 200 G = 100 a) Find out the equilibrium level of income. b) Suppose G increases to 125 what is the new equilibrium level of income? c) What level of G is needed to achieve a target income of 200?
1
Expert's answer
2020-05-25T09:49:40-0400

"Ye = C + I + G,"  

The equilibrium level of income becomes; Ye=50+0.80 (Y-T) +200+ 100

Ye=350+ 0.80(Y-T)

If G increases to 125 , then Ye=50+0.80 (Y-T) +200+ 125

Ye=375+ 0.80(Y-T)

If the income is 200, then G becomes

200=50+0.80 (Y-T) +200+ G

G=-50-0.80(Y+T)